# UniEmoX: Cross-modal Semantic-Guided Large-Scale Pretraining for Universal Scene Emotion Perception

### v2

@@ III-B Network Architecture
$\Upsilon=Swish\left(\sum_{i=1}^{\kappa}\left(softmax\left(\alpha\eta_{i}\right)\odot\alpha+softmax\left(\beta\xi_{i}\right)\odot\beta\right)\right)+\epsilon.$ (3)
For fusion strategy $\Gamma_{3}$, multiple sets of linear transformation matrices $\phi_{i}\in\mathbb{R}^{C_{3}\times C_{3}}$ and $\varphi_{i}\in\mathbb{R}^{C_{3}\times C_{3}}$ are defined for $\forall i\in\left\{1,2,\dots,\kappa\right\}$. This strategy primarily utilizes multi-head attention applied to two input vectors $\alpha,\beta\in\mathbb{R}^{N\times C_{3}}$, combined with a gating mechanism, to produce a refined output for feature fusion. This process can be represented as follows
$\Delta_{i}$ $=softmax\left(\alpha\phi_{i}\right)\odot\alpha+softmax\left(\beta\varphi_{i}\right)\odot\beta,$ (4a)
$\Delta_{con}$ $=concat(\Delta_{1},\Delta_{2},\cdots,\Delta_{\kappa}),$ (4b)
$\Delta_{m}$ $=ReLU\left(\Delta_{con}W_{1}+b_{1}\right)W_{2}+b_{2},$ (4c)
$\Delta_{g}$ $=\Delta_{m}\odot\sigma\left(\Delta_{m}W_{3}+b_{3}\right),$ (4d)
$\Upsilon$ $=ReLU\left(\Delta_{m}\odot\sigma\left(\Delta_{g}\right)\right)+\epsilon.$ (4e)
where $W_{1}\in\mathbb{R}^{\left(C_{3}\times H\right)\times C_{3}}$, $W_{2}\in\mathbb{R}^{C_{3}\times C_{3}}$, $b_{1},b_{2}\in\mathbb{R}^{C}$, $\sigma$ represents the Sigmoid activation function.
Distilling Rich Semantic Knowledge from CLIP. CLIP has demonstrated the potential for cross-modal learning between language and vision. Traditionally, language and vision processing are considered independent tasks. However, CLIP simultaneously processes both types of information, leveraging their contrastive relationship to significantly enhance the model’s understanding capabilities. Inspired by the success of CLIP, we adopt two approaches to develop a robust emotional pretraining framework: first, incorporating the successful vision-language contrastive pretraining paradigm, and second, efficiently distilling the visual semantic knowledge learned by CLIP into the emotional pretraining model. To achieve these goals, we implement the following specific measures:
We construct descriptive texts by naturally linking label attribute from the large-scale emotional dataset EmoSet, thereby creating numerous emotion image-text pairs. In incorporating the vision-language contrastive pretraining paradigm, we do not optimize contrastive loss as CLIP do. Given that CLIP’s training process utilized extensive data and hardware resources, its training performance is outstanding. To fully leverage CLIP’s pretrained model, we design a similarity-based contrastive loss function to optimize the training process. Specifically, we input descriptive text $T$ and full-scene image $I$ into CLIP’s text encoder and image encoder, respectively, obtaining text features $\Lambda\in\mathbb{R}^{N\times C_{3}}$ and visual features $\Omega\in\mathbb{R}^{N\times C_{3}}$. We first apply L2 normalization to the CLIP-processed text features $\Lambda$, visual features $\Omega$, and discriminative visual emotional representation $\Upsilon$. This process can be represented as follows
$\mathrm{X^{\prime}}=\left[\frac{\mathrm{X_{c}}}{\sqrt{\left\|\mathrm{X_{c}}\right\|^{2}}}\right]_{c=1}^{C_{3}}$ (5)
where $\mathrm{X}\in\{\Lambda,\Omega,\Upsilon\}$, $\Lambda$, $\Omega$, $\Upsilon$ is an $N\times C_{3}$ matrix, with each row vector $\Lambda_{c}$, $\Omega_{c}$, $\Upsilon_{c}$ representing the n-th row of matrix $\Lambda$, $\Omega$, $\Upsilon$, respectively.
Then, we calculate the sample correlation matrices between the visual emotional representation $\Upsilon^{\prime}$ and the CLIP-processed text features $\Lambda^{\prime}$, as well as between the CLIP-processed visual semantic representation $\Omega^{\prime}$ and the CLIP-processed text features $\Lambda^{\prime}$. Both sample correlation matrices are subjected to L2 normalization, resulting in matrices $\mathrm{A}\in\mathbb{R}^{N\times N}$ and $\mathrm{C}\in\mathbb{R}^{N\times N}$. This process can be represented as follows
$\mathrm{A}=\left[\frac{\left(\Upsilon^{\prime}\Lambda^{\prime T}\right)_{n}}{\sqrt{\left\|\left(\Upsilon^{\prime}\Lambda^{\prime T}\right)_{n}\right\|^{2}}}\right]_{n=1}^{N}\quad\mathrm{C}=\left[\frac{\left(\Omega^{\prime}\Lambda^{\prime T}\right)_{n}}{\sqrt{\left\|\left(\Omega^{\prime}\Lambda^{\prime T}\right)_{n}\right\|^{2}}}\right]_{n=1}^{N}.$ (6)
Finally, a similarity-based contrastive loss is employed for training to align our model’s vision-language contrastive similarity with that of CLIP. The ultimate goal is to transfer CLIP’s vision-language knowledge to our model. The similarity-based contrastive loss function $\mathcal{L}_{2}$ is expressed as follows
$\mathcal{L}_{2}=1-\frac{1}{\left|N\right|}{\textstyle\sum_{n\in N}}\frac{A_{n}\cdot C_{n}^{T}}{\sqrt{\left\|A_{n}\right\|^{2}}\sqrt{\left\|C^{T}_{n}\right\|^{2}}}.$ (7)
During the process of distilling the visual semantic knowledge learned by CLIP into the emotional pretraining model, we employ a visual feature similarity loss to transfer CLIP’s visual encoding capabilities to our model.
$\mathcal{L}_{3}=1-\frac{1}{\left|N\right|}{\textstyle\sum_{n\in N}}\frac{\Upsilon^{\prime}_{n}\cdot\Omega^{\prime T}_{n}}{\sqrt{\left\|\Upsilon^{\prime}_{n}\right\|^{2}}\sqrt{\left\|\Omega^{\prime T}_{n}\right\|^{2}}}.$ (8)
During the training phase, the overall optimization objective is defined as $\mathcal{L}_{t}=\mathcal{L}_{1}+\mathcal{L}_{2}+\mathcal{L}_{3}$.
III-C The Optimization Objectives in the Fine-Tuning Stage
@@
During the fine-tuning phase of downstream tasks, we employ mixup[37] and soft targets[38] to improve the model’s performance with small sample sizes, imbalanced distributions, or noisy data. Assuming the predicted value of the sample is $x\in\mathbb{R}^{N\times d}$ and the target label is $y\in\mathbb{R}^{d}$. The soft target cross entropy loss $\mathcal{L}_{f}$ can be expressed as follows
$\mathcal{L}_{f}=\frac{1}{N}\sum_{n=1}^{N}\left(-x_{n,y_{n}}+\log{\left(\sum_{i=1}^{d}\exp\left(x_{n,i}\right)\right)}\right).$ (9)
@@ IV-B Dataset Statistics
V Experiments
@@
Params (M) FLOPs (G) ViT-B ViT-L CNN UBE[10] FI[45] Emotion6[43] Emo8 (ours) CAER-S[46] HECO[44]
@@
Stimuli-awareTIP ’21 [4] - - ResNet-50, ResNet-101 Label - 72.42 61.62 - - -
MDANCVPR ’22 [22] 48.79 101.4 ResNet-101 Label - - 61.66 - - -
@@
UniEmoX 329.54 20.78 100 or 200 CLIP B/16 + Pixel 84.43 $\uparrow$ 0.6 69.33$\downarrow$ 3.09 65.67 $\uparrow$ 1.49 81.33 $\uparrow$ 0.55 92.09 $\uparrow$ 0.92 47.25$\downarrow$ 1.12
@@
Component Image Reconstruction Semantic Knowledge Psychological Prior SVEA PVEA
$\mathcal{L}_{1}$ $\mathcal{L}_{3}$ $\mathcal{L}_{2}$ UBE[10] FI[45] Emo8 (Ours) CAER-S[46] HECO[44]
$S_{1}$ 68.71 54.79 75.44 91.86 35.34
$S_{2}$ 69.61 58.69 73.78 83.07 33.50
$S_{3}$ 62.87 56.45 75.33 91.09 35.34
$S_{4}$ 53.89 56.20 74.22 91.98 35.03
$S_{5}$ 71.41 60.30 75.33 92.84 40.63
$S_{6}$ 74.70 61.30 77.78 92.98 40.94
$S_{7}$ 84.43 69.33 81.33 92.09 47.25
@@
UBE FI Emo8 (ours) CAER-S HECO
0.25 75.45 61.27 77.11 92.67 33.71
0.50 75.75 61.30 78.00 92.60 36.76
0.75 71.41 60.30 75.33 92.84 40.63
0.85 71.41 60.44 75.78 92.47 37.88
@@
UBE FI Emo8 (ours) CAER-S HECO
$\Gamma_{1}$ 76.80 61.45 77.22 90.88 35.74
$\Gamma_{2}$ 76.35 62.57 76.22 90.15 42.16
$\Gamma_{3}$ 76.80 62.36 76.56 89.51 37.37
$\Gamma_{4}$ 71.41 60.30 75.33 92.84 40.63
@@ V-A Datasets
By linking the emotional label attributes of the EmoSet[34] dataset through natural language logic, we construct 118,000 emotion text-image pairs for pretraining UniEmoX. To evaluate the effectiveness and generalization ability of UniEmoX, we conduct comparison experiments, ablation studies, and visual analyses on five public benchmark datasets: CAER-S[46], HECO[44], UBE[10], Emotion6 [43] and FI[45] as well as our Emo8 dataset. Due to space constraints, descriptions of downstream tasks and their corresponding datasets are provided in Appendix Section I.
@@ V-D Ablation Study
To verify the effectiveness of the various component combinations, we conduct ablation experiments. The specific steps are as follows: we conduct 150 epochs of self-supervised pre-training, experimenting with various combinations of components. (Note that the results from 400 epochs of self-supervised pre-training, where all components are combined, are referred to as $S_{6}$. The results from 200 epochs of self-supervised pre-training, following the loading of the pre-trained model with all components combined, are referred to as $S_{7}$.) and then fine-tune the pretrained models on downstream datasets. The experimental results are shown in TABLE IV. Based on these results, we can draw three conclusions:
@@
    By comparing $S_{1}$ with $S_{4}$, $S_{2}$ with $S_{4}$, and $S_{3}$ with $S_{4}$, we respectively validate the effectiveness of psychological theory prior knowledge, visual semantic knowledge distilled loss function $\mathcal{L}_{3}$ from CLIP, and reconstruction loss function $\mathcal{L}_{1}$ in some datasets;
@@
    By comparing $S_{4}$ with $S_{5}$, and $S_{4}$ with $S_{6}$, we confirm the effectiveness of the contrastive similarity loss function $\mathcal{L}_{2}$ across all datasets, which significantly improves recognition accuracy in downstream datasets;
@@
    By comparing $S_{5}$ with $S_{6}$, we find that increasing the number of self-supervised pretraining epochs can enhance performance in downstream datasets. By comparing $S_{7}$ with $S_{6}$, we find that initializing with a pre-trained model significantly enhances the performance of downstream datasets in environments with limited experimental resources.
To assess the impact of masking ratios and various fusion strategies on downstream emotion analysis tasks, we conduct ablation experiments. The specific steps are as follows: we perform 150 epochs of pre-training with the model using different masking ratios and fusion strategies, followed by fine-tuning the pre-trained model on downstream datasets. The experimental results are shown in TABLE. V and TABLE. VI. Based on these results, we can draw one conclusion:
@@
    Although MAE research suggests that models pre-trained with a 0.75 masking ratio perform better during fine-tuning, we find that models pre-trained with a 0.5 masking ratio learn emotional representations more effectively in most downstream datasets, including UBE, FI, and Emo8. Conversely, models pre-trained with a 0.75 masking ratio excel only in fine-tuning a few datasets, such as CAER-S and HECO.
  2. 2.
    $\Gamma_{1}$ and $\Gamma_{2}$ exhibited superior performance, with $\Gamma_{2}$ proving more effective for datasets characterized by significant semantic gaps. This is particularly relevant for HECO and FI, which possess relatively large data volumes and are not subjected to stringent data quality filtering. While $\Gamma_{3}$ utilizes additional parameters, its performance on downstream datasets is suboptimal. Conversely, $\Gamma_{4}$, as a baseline method with minimal parameters, underperform across most downstream datasets and is more suitable for addressing smaller semantic gaps.

### v3

@@ III-D1 Contrastive Feature Alignment Loss
  • Dual Normalization: By applying $\ell_{2}$-normalization: to each row of $A$ and $C$, the optimization objective shifts from absolute value matching to directional alignment. Compared to traditional knowledge distillation based on KL divergence, this design relaxes constraints to preserve intrinsic geometric properties of heterogeneous features.
  • Text Bridging: The shared text features $\Lambda^{\prime}$ establish a unified semantic reference frame, enabling indirect alignment of $\Upsilon^{\prime}$ and $\Omega^{\prime}$ while mitigating structural mismatch.
As presented in TABLE IV, incorporating CFA $\mathcal{L}_{2}$ Loss enhances emotion recognition accuracy by an average of 2.02% on downstream datasets compared to its absence.
III-D2 Semantic-aware Visual Alignment
To effectively distill the visual semantic knowledge encoded in CLIP’s feature representations ($\Omega_{n}\in\mathbb{R}^{C_{3}}$) into our emotion-aware pretraining framework, we propose a Semantic-aware Visual Alignment (SVA) Loss $\mathcal{L}_{3}$, is defined as follows:
$\mathcal{L}_{3}=1-\frac{1}{N}\sum_{n=1}^{N}\frac{\Upsilon_{n}\cdot\Omega_{n}}{\|\Upsilon_{n}\|_{2}\|\Omega_{n}\|_{2}}.$ (10)
This design facilitates geometric alignment between $\Upsilon_{n}$ and $\Omega_{n}$ in the feature space, while intentionally avoiding strict numerical equivalence. Our theoretical rationale stems from two perspectives:
  • Directional Semantic Preservation: The cosine similarity metric focuses exclusively on angular alignment rather than vector magnitude matching. This preserves the intrinsic geometry of emotion-specific representations, preventing over-constraint that could force $\Upsilon_{n}$ to collapse into generic visual semantics. Mathematically, the loss optimizes:
    $\min\theta(\Upsilon_{n},\Omega_{n})\quad\text{subject to}\quad\|\Upsilon_{n}\|_{2}\neq\|\Omega_{n}\|_{2}.$ (11)
    where $\theta$ denotes the inter-feature angle.
  • Hierarchical Knowledge Transfer: Complete vector alignment would reduce our model to replicating CLIP’s visual encoding behavior, thereby limiting its capacity to learn higher-order affective patterns. The magnitude-agnostic design creates a controllable knowledge distillation gap, allowing simultaneous acquisition of foundational visual semantics from CLIP and emotion-discriminative features through task-specific learning.
As presented in TABLE IV, incorporating SVA $\mathcal{L}_{3}$ Loss enhances emotion recognition accuracy by an average of 4.16% on downstream datasets compared to its absence.
III-E The Optimization Objectives in the Fine-Tuning Stage
@@
During the fine-tuning phase of downstream tasks, we employ soft targets[40] to improve the model’s performance. Assuming the predicted value of the sample is $x\in\mathbb{R}^{N\times d}$, the soft target cross entropy loss $\mathcal{L}_{f}$ can be defined as follows:
$\mathcal{L}_{\text{soft}}=-\frac{1}{N}\sum_{n=1}^{N}\sum_{i=1}^{d}t_{n,i}\log\left(\sigma(x_{n,i})\right),$ (12)
where $x_{n}\in\mathbb{R}^{d}$ is the output logits (un-normalized prediction) for the $n$-th sample, $t_{n,i}$ is the target soft label for the $n$-th sample and $i$-th class (i.e., the target probability for each class), $\sigma(x)$ is the softmax function applied to the logits, producing the predicted class probabilities.
The purpose of this design is to maximize the probability of the correct class while minimizing the probabilities of other classes, enabling the model to learn more accurate classification capabilities.
@@
IV Emo8 Dataset
IV-A Dataset Collection
IV-B Dataset Statistics
@@
Params (M) FLOPs (G) ViT-B ViT-L CNN UBE[10] FI[46] Emotion6[44] EmoSet[38] Emo8 (ours) CAER-S[47] HECO[45]
@@
Stimuli-awareTIP ’21 [4] - - ResNet-50, ResNet-101 Label - 72.42 61.62 78.40 - - -
MDANCVPR ’22 [24] 48.79 101.4 ResNet-101 Label - - 61.66 75.75 - - -
@@
UniEmoX 329.54 20.78 100 or 200 CLIP B/16 + Pixel 84.43 $\uparrow$ 0.6 69.33$\downarrow$ 3.09 65.67 $\uparrow$ 1.49 78.71 $\downarrow$ 1.48 81.33 $\uparrow$ 0.55 92.09 $\uparrow$ 0.92 47.25$\downarrow$ 1.12
@@
Component Image Reconstruction Semantic Knowledge Psychological Prior
w/o $\mathcal{L}_{1}$ w/ $\mathcal{L}_{1}$ $\Delta$ w/o $\mathcal{L}_{2}$ w/ $\mathcal{L}_{2}$ $\Delta$ w/o $\mathcal{L}_{3}$ w/ $\mathcal{L}_{3}$ $\Delta$ w/o fusion w/ fusion $\Delta$
UBE 83.38 84.43 $\uparrow$ 1.05 84.58 84.43 $\downarrow$ 0.15 82.64 84.43 $\uparrow$ 1.79 82.19 84.43 $\uparrow$ 2.24
FI 67.99 69.33 $\uparrow$ 1.34 68.82 69.33 $\uparrow$ 0.51 66.81 69.33 $\uparrow$ 2.52 69.56 69.33 $\downarrow$ 0.23
Emotion6 64.68 65.67 $\uparrow$ 0.99 63.18 65.67 $\uparrow$ 2.49 60.44 65.67 $\uparrow$ 5.23 59.45 65.67 $\uparrow$ 6.22
EmoSet 77.75 78.71 $\uparrow$ 0.96 77.89 78.71 $\uparrow$ 0.82 75.77 78.71 $\uparrow$ 2.94 78.63 78.71 $\uparrow$ 0.08
Emo8 80.44 81.33 $\uparrow$ 0.89 80.44 81.33 $\uparrow$ 0.89 78.11 81.33 $\uparrow$ 3.22 80.00 81.33 $\uparrow$ 1.33
CAER-S 92.68 92.09 $\downarrow$ 0.59 92.69 92.09 $\downarrow$ 0.60 90.39 92.09 $\uparrow$ 1.70 92.45 92.09 $\downarrow$ 0.36
HECO 46.23 47.25 $\uparrow$ 1.02 37.07 47.25 $\uparrow$ 10.18 35.54 47.25 $\uparrow$ 11.71 38.70 47.25 $\uparrow$ 8.55
Average $\uparrow$ 0.95 $\uparrow$ 2.02 $\uparrow$ 4.16 $\uparrow$ 2.69
@@
UBE FI Emotion6 EmoSet Emo8 (ours) CAER-S HECO
0.25 75.45 61.27 56.22 73.58 77.11 92.67 33.71
0.50 75.75 61.30 59.45 73.85 78.00 92.60 36.76
0.75 71.41 60.30 53.98 72.43 75.33 92.84 40.63
0.85 71.41 60.44 48.01 70.61 75.78 92.47 37.88
@@
UBE FI Emotion6 EmoSet Emo8 (ours) CAER-S HECO
$\Gamma_{1}$ 76.80 61.45 54.98 72.14 77.22 90.88 35.74
$\Gamma_{2}$ 76.35 62.57 54.98 72.73 76.22 90.15 42.16
$\Gamma_{3}$ 76.80 62.36 54.23 72.78 76.56 89.51 37.37
$\Gamma_{4}$ 71.41 60.30 53.98 72.43 75.33 92.84 40.63
TABLE VII: “w/ labels” indicates that the text descriptions include manually annotated emotional information during pretraining, whereas “w/o labels” indicates the absence of such annotations.
SVEA PVEA
UBE FI Emotion6 EmoSet Emo8 (ours) CAER-S HECO
w/ labels 84.43 69.33 65.67 78.71 81.33 92.09 47.25
w/o labels 85.03 67.67 65.67 77.98 80.78 92.34 45.62
Difference $\uparrow$ 0.6 $\downarrow$ 1.66 - $\downarrow$ 0.73 $\downarrow$ 0.55 $\uparrow$ 0.25 $\downarrow$ 1.63
V Experiments
@@ V-A Datasets
By linking the emotional label attributes of the EmoSet[38] dataset through natural language logic, we construct 118,000 emotion text-image pairs for pretraining UniEmoX. To evaluate the effectiveness and generalization ability of UniEmoX, we conduct comparison experiments, ablation studies, and visual analyses on five public benchmark datasets: CAER-S[47], HECO[45], UBE[10], Emotion6 [44], EmoSet and FI[46] as well as our Emo8 dataset. Due to space constraints, descriptions of downstream tasks and their corresponding datasets are provided in Appendix Section I.
@@ V-D Ablation Study
To verify the effectiveness of the various component combinations, we conduct ablation experiments. The specific steps are as follows: we initialize with a pre-trained model and conduct 200 epochs of multimodal pre-training with a diverse set of components, followed by fine-tuning on multiple downstream datasets. The experimental results are shown in TABLE IV. Based on these results, we can draw three conclusions:
@@
    By comparing w/o $\mathcal{L}_{1}$ vs. w/ $\mathcal{L}_{1}$, w/o $\mathcal{L}_{2}$ vs. w/ $\mathcal{L}_{2}$, w/o $\mathcal{L}_{3}$ vs. w/ $\mathcal{L}_{3}$, and w/o fusion vs. w/ fusion, we systematically evaluate the contributions of Image Reconstruction, Semantic Knowledge, and Psychological Prior to the model’s performance. Notably, the Semantic-aware Visual Alignment (SVA) Loss $\mathcal{L}_{3}$ plays the most critical role within the module combination.
@@
    Although certain component ($\mathcal{L}_{1}$, $\mathcal{L}_{2}$, and fusion) exhibit performance degradation on specific datasets such as CAER-S, UBE, and FI, their overall performance demonstrates an upward trend across all datasets on average. This result provides strong evidence supporting the stability of the proposed components.
To assess the impact of masking ratios and various fusion strategies on visual downstream emotion analysis tasks, we conduct ablation experiments. The specific steps are as follows: the model is first randomly initialized, followed by 150 epochs of pre-training using different masking ratios and fusion strategies, and then fine-tuning the pre-trained model on downstream datasets. The experimental results are shown in TABLE. V and TABLE. VI. Based on these results, we can draw one conclusion:
  1. 1.
    Although MAE research suggests that models pre-trained with a 0.75 masking ratio perform better during fine-tuning, we find that models pre-trained with a 0.5 masking ratio learn emotional representations more effectively in SVEA datasets. Conversely, models pre-trained with a 0.75 masking ratio excel in fine-tuning PVEA datasets.
  2. 2.
    $\Gamma_{1}$ and $\Gamma_{2}$ exhibited superior performance, with $\Gamma_{2}$ proving more effective for datasets characterized by significant semantic gaps. This is particularly relevant for HECO and FI, which possess relatively large data volumes and are not subjected to stringent data quality filtering.
@@
    While $\Gamma_{3}$ utilizes additional parameters, its performance on downstream datasets is suboptimal. Therefore, $\Gamma_{2}$ serves as a robust fusion method, more adaptable to datasets with varying qualities, cues, and semantic gaps.
  4. 4.
    Conversely, $\Gamma_{4}$, as a baseline method with minimal parameters, underperform across most downstream datasets and is more suitable for addressing smaller semantic gaps. More suitable for datasets like CAER-S, where image content is relatively simple and semantic cues are limited.
  5. 5.
    Overall, $\Gamma_{2}$ and $\Gamma_{3}$ outperform $\Gamma_{1}$ and $\Gamma_{4}$, as they incorporate additional parameters that enable the learning of more patterns, thereby providing stable performance across various data quality gaps. While $\Gamma_{4}$ and $\Gamma_{1}$ perform well on certain datasets, they tend to suffer from performance fluctuations when faced with datasets characterized by more complex semantic gaps, such as HECO and UBE.
The text descriptions corresponding to emotional images are generated by logically linking multiple emotion-related attribute labels using natural language. These attributes include brightness, color saturation, scene type, object type, facial expressions, action type, and overall emotional expression. Among them, brightness, color saturation, scene type, object type, facial expressions, and action type are intrinsic image features, present at both the content and structural levels. These attributes are predicted by the model rather than manually annotated. Therefore, when these text descriptions—constructed through natural language connections—are used for UniEmoX pretraining, no additional supervision signals are introduced, and no semantic knowledge is leaked. As a result, the entire process can be regarded as fully self-supervised learning. However, in our previous experiments, the text descriptions included overall emotional expression, which may have introduced some degree of semantic information. To further assess its impact on model performance, we conducted an ablation experiments. The experimental results are shown in TABLE. VII. Based on these results, we can draw one conclusion:
@@
    The results show that the impact varies across the six datasets, with the largest performance difference being approximately two percentage points. We hypothesize that the limited impact is due to the relatively small proportion of overall emotional expression within the complete natural language descriptions, which may not significantly interfere with the model’s learning process.